Graph y=5x.
NOTE- the x after 5 is an exponent.
Thanks,
Pat
: Graph y=5x.
NOTE- the x after 5 is an exponent.
Thanks,
Pat

Answer by jim_thompson5910(9165)

I've noticed that you've posted this before. Is there something that isn't making any sense? Well here's the original solution.


Table of Contents:

Table
Graph

In order to graph y=5^x, we need to plot a few points.


Let's find the y value when x=-1 note: you can start at any x value.


y=5^x Start with the given equation.


y=5^-1 Plug in x=-1.


y=0.2 Simplify.


So when x=-1, y=0.2. So we have the point (-1,0.2).


----------------------------


Let's find the y value when x=0


y=5^x Start with the given equation.


y=5^0 Plug in x=0.


y=1 Simplify.


So when x=0, y=1. So we have the point (0,1).


----------------------------


Let's find the y value when x=1


y=5^x Start with the given equation.


y=5^1 Plug in x=1.


y=5 Simplify.


So when x=1, y=5. So we have the point (1,5).


----------------------------


Let's find the y value when x=2


y=5^x Start with the given equation.


y=5^2 Plug in x=2.


y=25 Simplify.


So when x=2, y=25. So we have the point (2,25).


----------------------------


Now let's make a table of the values we just found.





Jump to Top of Page

Table of Values:


TEST: -1
xy
-10.2
01
15
225

Now let's plot the points:



Photobucket - Video and Image Hosting




Jump to Top of Page

Graph:


Now draw a curve through all of the points to graph y=5^x:


Photobucket - Video and Image Hosting Graph of y=5^x